(21) Water has a mass per mole of 18.0 g/mol, and each water molecule (H2O) has 10 electrons.

(a) How many electrons are there in 1.0 L of water?
(b) What is the net charge of all these electrons?

(22) Two point charges are fixed on the y-axis: a negative point charge

q1 = ?25.0 µC at y1 = 0.22 m and a positive charge q2 at y2 = 0.34 m. A third point charge q = 8.4 µC

is fixed at the origin. The net electrostatic force exerted on the charge q by the other two charges has a magnitude 27.0 N and points in the +y-direction.

(21) Water has a mass per mole of 18.0 g/mol,

(23) Determine the magnitude of

q2.
Answer #1

21

number of electrons in one molecule = 10

number of molecules in 1 mole = 6.023 x 1023

number of electrons in 1 mole = 6.023 x 1024

1 L of water = 1000 g of water

number of moles in 1 L water = 1000 /18 = 55.6

Total number of electrons = (6.023 x 1024) (55.6) = 3.35 x 1026

b)

net charge = q = (3.35 x 1026) (1.6 x 10-19) = 5.36 x 107 C